OpenStudy (anonymous):

Calculate the equation of the line between the points (-6, -1) and (2, -1)

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

is this the equation?

OpenStudy (anonymous):

for the slope u have to calculate tht first

OpenStudy (anonymous):

so apparently slope =0

OpenStudy (anonymous):

y=ax+b (a=0) hence y=b

OpenStudy (anonymous):

the answers are either a.y=-1 b.x=-1 c.y=-8x-1 d.y=-x-8

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

thanx hahd

OpenStudy (anonymous):

np